# Differentiable Weightless Neural Networks

### v2

@@ 4.1 DWNs on FPGAs
Table 1: Implementation results for DWNs and prior efficient inference models on a Z-7045 FPGA with input buffers and decompression implemented, with data being passed to the board using its I/O pins, which have a bandwidth of 112 bits per cycle, and the clock is limited to 200 MHz for a fair comparison to prior work. For results in out-of-context mode, where data is assumed to already be on the board and the clock is set to 250 MHz, see Appendix D. Model could not be synthesized; hardware values are approximate. ULEEN used augmented data for MNIST; we present MNIST results with and without augmentation.
@@
Table 2: FPGA implementation results comparing DWNs and other LUT-based Neural Networks. LogicNets, PolyLUT and NeuralLUT results are taken from their respective papers. New LogicNets results taken from their official GitHub (Umuroglu et al., 2020b)
Dataset Model Accuracy LUT FF DSP BRAM Fmax (MHz) Latency (ns) $\bm{\times}$Delay (LUT×ns)
MNIST PolyLUT 96% 70673 4681 0 0 378 16 1.1e+06
NeuraLUT 96% 54798 3757 0 0 431 12 6.6e+05
DWN ($n$=6; sm) 97.8% 2092 1757 0 0 873 5.8 1.2e+04
JSC ($<$ 75%) LogicNets 71.8% 37931 810 0 0 427 13 4.9e+05
LogicNets (sm)* 69.8% 244 270 0 0 1353 5 1.2e+03
LogicNets (lg)* 73.1% 36415 2790 0 0 390 6 2.2e+05
PolyLUT 72% 12436 773 0 0 646 5 6.2e+04
NeuraLUT 72% 4684 341 0 0 727 3 1.4e+04
DWN ($n$=6; sm) 73.7% 134 106 0 0 1361 1.5 2.0e+02
JSC ($75\%-76\%$) PolyLUT 75% 236541 2775 0 0 235 21 5.0e+06
NeuraLUT 75% 92357 4885 0 0 368 14 1.3e+06
DWN ($n$=6; md) 75.6% 2144 1457 0 0 903 5.6 1.2e+04
JSC ($\geq$ 76%) hls4ml 76.2% 63251 4394 38 0 200 45 2.8e+06
DWN ($n$=6; lg) 76.3% 6302 4128 0 0 695 10.1 6.4e+04
@@ Appendix D Out-of-Context Results
Dataset Model Test AccuracyParameter Size (KiB) Time/Sample (ns) Throughput (Samples/s) LUTs (1000s)
MNIST DWN ($n$=2; lg) 98.27 5.9 4 250 M 6.5
DWN ($n$=6; sm) 97.80 11.7 4 250 M 2.1
DWN ($n$=6; lg) 98.31 23.4 4 250 M 4.1
DWN ($n$=6; lg; +aug) 98.77 23.4 4 250 M 4.1
FashionMNIST DWN ($n$=2) 89.12 7.8 4 250 M 8.1
DWN ($n$=6) 89.01 31.3 4 250 M 6.2
KWS DWN ($n$=2) 70.92 2.9 4 250 M 3.3
DWN ($n$=6) 71.52 12.5 4 250 M 3.3
ToyADMOS/car DWN ($n$=2; sm) 86.68 0.9 4 250 M 1.0
DWN ($n$=6; sm) 86.93 3.1 4 250 M 0.8
DWN ($n$=6; lg) 89.03 28.1 4 250 M 5.5
CIFAR-10 DWN ($n$=2) 57.51 23.4 4 250 M 45.7
DWN ($n$=6) 57.42 62.5 4 250 M 16.7

### v3

@@ 4.1 DWNs on FPGAs
Table 1: Implementation results for DWNs and prior efficient inference models on a Z-7045 FPGA with input buffers and decompression implemented, with data being passed to the board using its I/O pins, which have a bandwidth of 112 bits per cycle, and the clock is limited to 200 MHz for a fair comparison to prior work. For results in out-of-context mode and with the clock is set to 250 MHz, see Appendix D. Model could not be synthesized; hardware values are approximate. ULEEN used augmented data for MNIST; we present MNIST results with and without augmentation.
@@
Table 2: FPGA implementation results comparing DWNs with other LUT-based neural networks. Results for LogicNets, PolyLUT, and NeuraLUT are sourced from their respective papers, while new LogicNets results () are from their official GitHub (Umuroglu et al., 2020b). ”Time/Sample” denotes the inference time per sample (calculated as 1/Throughput). Latency refers to the total end-to-end time of the model execution.
Dataset Model Accuracy LUT FF DSP BRAM Fmax (MHz) Time/Sample Latency $\bm{\times}$Lat. (LUT×ns)
MNIST PolyLUT 96% 70673 4681 0 0 378 2.6ns 16.0ns 1.1e+06
NeuraLUT 96% 54798 3757 0 0 431 2.3ns 12.0ns 6.6e+05
DWN ($n$=6; sm) 97.1% 692 422 0 0 827 1.2ns 2.4ns 1.6e+03
DWN ($n$=6; md) 97.8% 2055 1675 0 0 873 1.1ns 4.6ns 9.4e+03
DWN ($n$=6; md) 97.9% 1413 1143 0 0 827 1.2ns 3.6ns 5.1e+03
DWN ($n$=6; lg) 98.3% 4082 3385 0 0 827 1.2ns 6.0ns 2.4e+04
JSC LogicNets 71.8% 37931 810 0 0 427 2.3ns 13.0ns 4.9e+05
LogicNets (sm)* 69.8% 244 270 0 0 1353 0.7ns 5.0ns 1.2e+03
LogicNets (lg)* 73.1% 36415 2790 0 0 390 2.6ns 6.0ns 2.2e+05
PolyLUT 72% 12436 773 0 0 646 1.5ns 5.0ns 6.2e+04
NeuraLUT 72% 4684 341 0 0 727 1.4ns 3.0ns 1.4e+04
DWN ($n$=6; sm) 71.1% 20 22 0 0 3030 0.3ns 0.6ns 1.3e+01
DWN ($n$=6; sm) 74.0% 110 72 0 0 1094 0.9ns 1.5ns 2.0e+02
JSC PolyLUT 75% 236541 2775 0 0 235 4.3ns 21.0ns 5.0e+06
NeuraLUT 75% 92357 4885 0 0 368 2.7ns 14.0ns 1.3e+06
DWN ($n$=6; md) 75.6% 720 457 0 0 827 1.2ns 3.6ns 2.6e+03
JSC hls4ml 76.2% 63251 4394 38 0 200 - 45.0ns 2.8e+06
DWN ($n$=6; lg) 76.3% 4972 3305 0 0 827 1.2ns 7.3ns 3.6e+04
@@ Appendix D Out-of-Context Results
Dataset Model Test AccuracyParameter Size (KiB) Time/Sample Throughput (Samples/s) LUTs (1000s)
MNIST DWN ($n$=2; lg) 98.27 5.9 4ns 250 M 6.5
DWN ($n$=6; sm) 97.80 11.7 4ns 250 M 2.1
DWN ($n$=6; lg) 98.31 23.4 4ns 250 M 4.1
DWN ($n$=6; lg; +aug) 98.77 23.4 4ns 250 M 4.1
FashionMNIST DWN ($n$=2) 89.12 7.8 4ns 250 M 8.1
DWN ($n$=6) 89.01 31.3 4ns 250 M 6.2
KWS DWN ($n$=2) 70.92 2.9 4ns 250 M 3.3
DWN ($n$=6) 71.52 12.5 4ns 250 M 3.3
ToyADMOS/car DWN ($n$=2; sm) 86.68 0.9 4ns 250 M 1.0
DWN ($n$=6; sm) 86.93 3.1 4ns 250 M 0.8
DWN ($n$=6; lg) 89.03 28.1 4ns 250 M 5.5
CIFAR-10 DWN ($n$=2) 57.51 23.4 4ns 250 M 45.7
DWN ($n$=6) 57.42 62.5 4ns 250 M 16.7